Suppose I am asked to write a 50-page document. I have looked at the requirements and determine that I can write five pages per day for 10 days. I explain to the customer that I earn $10 per hour, and we agree that I will work 10 hours per day. However, I am paid for the number of hours I work and I report the number of pages I write each day, the hours worked, and the cost.

Day #1--> Worked 10 hours; wrote 5 pages. The customer accepted the pages and paid me the agreed upon amount for a day's work ($100). I am clearly on budget and on schedule.

Day #2--> I again worked 10 hours, and produced the planned 5 pages. The customer paid me $100. I am still on budget and on schedule.

Day #3--> I ran into a problem, and only delivered 2 pages. Also, I worked an extra 2 hours trying to fix the problem (I worked 12 hours that day). Since I get paid $10/hour, I submit a bill to the customer for 12 hours ($120).

(a) At this rate, how much will this project cost (EAC)

(b) At this rate, how much should the customer expect to spend (ETC)
Answer #1

Given: No. of pages to complete = 50

Duration of work = 10 days

No. of pages per day = 5

No. of hours per day = 10

Cost per hour = $10

Actual cost for project (if everything is on schedule) = 10*10 = $100 per day

Total cost = $100*10 = $1000 for the project

However, day 1 and day 2 are as per schedule.

Day 3- Work completed = 2 out of 5 and hours = 12

Actual work that is to be completed after 3 days = 15 pages

Actual budget for 3 days = $300

Work completed after 3 days = 12 pages

Percentage completion of work after 3 days = (12/15)*100 = 80%

Planned value for 80% = $300*80% = $240

Budget spent for 3 days (Earned value) = $320

Currently, the project is over-budgeted

CPI = PV/EV = 240/320 = 0.75

Original budget = $1000

  1. At this rate, how much will the project cost? (EAC)

Ans: EAC is given by formula BAC/CPI

EAC = 1000/0.75 = 1333.33

At this rate, the project will cost $1333.33

  1. AT this rate, how much should the customer expect to spend? ETC

Ans: ETC is given by formula = EAC-AC

AC is amount that is already spent for the project

AC = $100+$100+$120 = $320

ETC = $1333.33- $320 = $1013.33

At this rate, the customer is expected to spend $1013.33 for completion of project.